A flowerpot falls from a window sill 33.9 m above the sidewalk. What is the velocity of the flowerpot when it strikes the ground? The acceleration of gravity is 9.81 m/s 2 .


1
Expert's answer
2020-09-28T08:08:30-0400
"v^2=2gh\\v^2=2(9.81)(33.9)\\v=25.8\frac{m}{s}"